kekeman:

Given vector u = 30 ⟨cos 60°, sin 60°⟩, what are the magnitude and direction of −6u? A. 210; 60° B. 210; 240° C. 180; 240° D. -180; 60°

OLIVER69:

Your answer is D because 60 + 60 = 120 then add 30 to 120, which equals 180, then you know it has to be 60 degrees because cos60+sin60, which you can solve on a calculator. Makes sense?

kekeman:

@oliver69 wrote:
Your answer is D because 60 + 60 = 120 then add 30 to 120, which equals 180, then you know it has to be 60 degrees because cos60+sin60, which you can solve on a calculator. Makes sense?
Okay so it would be -180; 60°?

OLIVER69:

@kekeman wrote:
@oliver69 wrote:
Your answer is D because 60 + 60 = 120 then add 30 to 120, which equals 180, then you know it has to be 60 degrees because cos60+sin60, which you can solve on a calculator. Makes sense?
Okay so it would be -180; 60°?
yes I do believe it would be

kekeman:

Thank you

Narad:

\ \[-6u=180\left( \frac{- 1 }{ 2 },\frac{- \sqrt{3} }{ 2 } \right)\] \[\left| -6u \right|=180\] \[\theta=180+60=240\]

kekeman:

Ahhhhhhhhhhhhhhhhhhhhh I got it wrong damn it